# Non-overlapping, Schwarz-type Domain Decomposition Method for Physics and Equality Constrained Artificial Neural Networks

### v1

@@ 4.1.3 Helmholtz equation with a multi-scale solution
The Helmholtz equation governs propagation phenomena. Its solution on partitioned domains requires specialized treatment due to its unique characteristics. Unlike the Poisson’s equation, the Helmholtz equation’s discretization typically produces a symmetric but non-positive definite matrix [14]. Transmission conditions that work well for the Laplace’s equation does not readily extend to handle the Helmholtz equation. Separate domain decomposition methods have been proposed to tackle Poisson’s and Helmholtz equations (e.g. [28]). To this end, solving the Helmholtz and Poisson’s equations with a unified domain decomposition framework presents substantial technical challenges.
@@
We use the same uniform global mesh of $160\times 160$ as adopted in [24]. We partitioned our domain with a $5\times 5$ decomposition. In our Algorithm 1, the penalty scaling factor for BC constraints, $\eta_{B}$, is set to 10, and the smoothing constant $\zeta$ is set to 0.999. Our neural network and optimizer configurations remain identical to those we have used in the Poisson’s equation investigations. The exact solution to this problem is not known. Therefore, to evaluate the quality of the predictions, the same finite difference method, as adopted in [24], on a mesh of $360\times 360$ is used to obtain the reference solution.
Figure 7: Multi-scale Helmholtz equation: (a) finite difference solution on a $5\times 5$ partitioned domain, (b) predicted solution, (c) absolute point-wise error.
@@
Figure 8: Multi-scale Helmholtz equation: Evolution of (a) PDE constraints, (b) BC constraints, (c) objective functions, (d) relative $\mathit{l^{2}}$ norms, (e) the ratio $\beta/\alpha$, (e) the ratio $\gamma/\alpha$ for six specific subdomains.
To illustrate the performance of the method, we select six specific subdomains based on the symmetry of the solution, and track the evolution of the loss terms, the relative $\mathit{l^{2}}$ norms, and the interface parameter ratios, as shown in Figure 8. At the early stages of training, most subdomains, except the central one (2,2), fall into local minima with low values for PDE, boundary constraints and objective functions, as shown in Figures 8(a-c). This behavior arises from the challenging problem setup: the source term, with its narrow width and high magnitude, is concentrated in the central subdomain, while the surrounding subdomains have near-zero source terms. The high wave number further complicates the reduction of the PDE constraint in the central subdomain. As the PDE constraint in this region decreases to around $10^{-3}$, its solution improves, gradually propagating information to neighboring subdomains. This leads to an increase in the objective functions (interface losses) in adjacent subdomains, pulling them out of their local minima, as reflected in the noticeable rise in constraint values. The interplay between satisfying the boundary conditions and improving the physics constraints drives the reduction of the relative $\mathit{l^{2}}$ norms, as seen in Figure 8(d). This process continues over 30,000 iterations towards convergence.
@@ 4.2.1 Inferring thermal conductivity in multilayered materials
The technical challenge lies in inferring two independent $\kappa$ values from sparse data. We generate synthetic measurement data at ten locations using pseudo sensors that are equidistantly positioned. These sensors are placed between $y=0.2$ and $y=0.8$, situated 0.05 units away from the interface. Using a $2\times 1$ domain decomposition, each subdomain is assigned to a distinct layer. The global domain includes $64\times 64$ randomly distributed collocation points, comprising residual, boundary, and interface points, as shown in Figure 9. For evaluation, we employ a uniform $128\times 128$ mesh. The same neural network architecture, hyperparameters, and optimizer settings from the forward Poisson’s problem, as shown in Figure 2, are used here. It is important to note that the additional data constraint shares the same default penalty scaling factor, $\eta_{B}=1$, as the boundary condition constraint. The unknown conductivity of each layer is treated as a model parameter to be learned, i.e., $\kappa_{k}\in\theta_{k}$ for $k\in\{1,2\}$, with initial values randomly assigned between 0 and 1.
@@
Figure 11 depicts the evolution of two accuracy metrics, constraints, objective functions, and interface parameters, as well as their ratios for both layers. The relative $\mathit{l^{2}}$ norms of the temperature and the absolute errors of two predicted conductivities, $\mathit{l^{\infty}}$, are shown in Figs. 11(a-b), respectively. During the early stages of training, both metrics exhibit rapid convergence, with the errors in $\mathcal{E}_{r}^{u}$ and $\mathcal{E}_{\infty}^{\kappa}$, separately decreasing by approximately five and three orders of magnitude. After a brief stabilization phase, further reductions occur around 700 outer iterations, coinciding with a slight decrease in the boundary condition and data constraints, in Fig. 11(c), and a notable rise in the objective functions, in Fig. 11(d). The increase in objective function values is attributed to the growth of interface parameters, seen in Fig. 11(e). As discussed in Section 3.2, increasing $q^{j}$ helps maintains a strong gradient when the corresponding operator $\mathcal{O}_{i}^{j}$ converges to small values, thereby enhancing convergence rate. Eventually, the objective functions stabilize, and both accuracy metrics reach convergence after approximately 1200 outer iterations. Additionally, Fig. 11(f) shows that the interface ratios also stabilize between 1 and 2.
@@ 4.2.2 Inferring thermal conductivity in functionally graded materials
Here, we adopt a similar problem used by Shukla et al. [22] to evaluate the XPINN approach. Instead of the map of the United States of America, we consider a butterfly-shaped domain $\Omega$. The geometry of the domain is defined as:
@@
Figure 13 displays the predicted distributions of temperature and thermal conductivity, along with their absolute errors. In Fig. 13(b), the temperature prediction errors are relatively uniform across the four subdomains, generally within the order of $10^{-3}$ and are primarily localized near the boundaries. In contrast, Fig. 13(d) reveals that the errors in conductivity predictions are mostly concentrated in the top subdomains, particularly in $(0,0)$, where the maximum error is $1.033\times 10^{-1}$. Comparing these errors with the true ranges shown in Figs. 13(a) and 13(c), both predictions demonstrate high accuracy, with temperature predictions being more precise than those for conductivity.
Figure 14: Inverse identification of spatially varying conductivity in functionally graded materials: Evolution of (a) relative $\mathit{l^{2}}$ norms for temperature, (b) relative $\mathit{l^{2}}$ norms for conductivity, (c) objective functions, (d) PDE constraints, (e) BC constraints, (f) data constraints, (g) interface parameter ratio $\alpha^{(k)}/\alpha^{(u)}$, (h) $\beta/\alpha^{(u)}$, (i) $\gamma/\alpha^{(u)}$ for all the four subdomains.
Figure 14 shows the evolution of the relative $\mathit{l^{2}}$ norms for temperature and conductivity, the objective functions, three constraints, and the interface parameter ratios across all subdomains. As seen in Fig. 14(a), $\mathcal{E}_{r}^{u}$ converges first, oscillating around $10^{-4}$ after approximately 2500 outer iterations, while $\mathcal{E}_{r}^{\kappa}$, in Fig. 14(b), continues to decrease gradually. Around the same outer iterations, the objective functions, which initially rise, begin to decrease rapidly in Fig. 14(c). The PDE, BC and data constraints, in Figs. 14(d-f), converge to approximately $10^{-6}$, $10^{-6}$, and $10^{-5}$, respectively. Among these, subdomain (1,0) exhibits a slower decay in $\mathcal{C}_{B}$, but ultimately achieves the lowest $\mathcal{E}_{r}^{\kappa}$ over the iterations. Due to the inclusion of the Dirichlet interface operator for $\kappa$, Fig. 14(g) shows the ratios of the interface parameters $\alpha^{(\kappa)}$ to $\alpha^{(u)}$. $\dfrac{\alpha^{(\kappa)}}{\alpha^{(u)}}$, as well as the other interface parameter ratios, $\dfrac{\beta}{\alpha^{(u)}}$ and $\dfrac{\gamma}{\alpha^{(u)}}$ in Figs. 14(h-i), fluctuate between 1.2 and 1.5 after an initial spike during the early stages of training. As noted earlier, the convergence of the relative $\mathit{l^{2}}$ norms does not necessarily require the convergence of the interface parameter ratios.
@@ 4.3 Parallel Performance
We assess the parallel efficiency of our domain decomposition method, using the same Poisson’s equation problem as outlined in Section 4.1.1. Using the Adam optimizer, we evaluate both weak and strong scaling properties, varying the number of subdomains from 2 to 32. These properties are derived from the average time costs recorded over 5 independent trials for each configuration. The numerical experiments are conducted on an Intel Xeon Platinum 8462Y+ architecture. Each subdomain is managed by an MPI process (mpi4py 3.1.5 package) bound to a single CPU core. A Cartesian communicator is created via Create_cart function and the interface information is sent and received via Isend and Irecv functions, respectively. It is important to note that our parallel implementation is straightforward; more advanced techniques such as overlapping communication with computation to improve scaling across many processors are not explored in this study.
@@ 4.3.1 Weak Scaling Analysis
Figure 15(a) captures the runtime for five outer iterations, segmented into computation and communication times. The total runtime slightly increase, approximately 60 seconds, across various domain decompositions, with computation time significantly exceeding communication time. This predominance is attributed to the unique design of our domain decomposition training procedure. Unlike other approaches (e.g. XPINN[20, 21]) where communication occurs every epoch, in our method, communication is scheduled for at least 50 inner epochs, which is possible because our formulation aims to infer the parameters of a generalized interface condition.
@@ 5 Conclusion
Domain decomposition methods are pivotal for advancing physics-informed or constrained machine learning techniques to tackle large-scale partial differential equation (PDE) challenges in a distributed setting. Our study introduces a novel Schwarz-type, non-overlapping domain decomposition method featuring a generalized interface condition. This method leverages physics-informed neural networks across distinct subdomains, efficiently solving both forward and inverse PDE problems. A significant innovation of our approach is its ability to uniformly address both Laplace’s and Helmholtz equations within the same framework, thereby obviating the need for ad-hoc tuning of neural network models. This unified framework contrasts sharply with traditional methods, which typically require separate domain decomposition strategies for each equation type. However, we should note, despite the common framework, each objective function will be unique to the PDE problem at hand because of the PDE-constrained nature of identifying the parameters of the interface condition.
Our proposed DDM utilizes the physics and equality constrained artificial neural networks (PECANN) framework [25] to formulate a constrained optimization problem for each subdomain. Unlike the original PECANN approach without a DDM, where boundary conditions solely constrain the PDE, in the present DDM, the boundary conditions and the PDE jointly constrain a generalized interface loss term, which is formulated as a linear combination of Dirichlet, Neumann, and tangential derivative continuity operators with parameters to be optimized. We employ an augmented Lagrangian method with a conditionally adaptive strategy that is unique to the present study to recast each constrained optimization problem into a dual unconstrained problem. Subdomains are independently trained using designated neural networks for a fixed number of epochs before synchronizing with neighboring subdomains, thus minimizing communication overhead while promoting the independent learning of interface parameters. This strategy significantly boosts both the efficiency and the effectiveness of the method.
We have rigorously demonstrated the robust parallel performance and versatility of our method on a variety of forward and inverse problems using complex domain partitioning strategies. Employing the MPI model, both weak and strong scaling analyses have indicated solid performance up to 32 processes.

### v2

@@ 4.1.3 Helmholtz equation with a complex-valued solution
The Helmholtz equation governs propagation phenomena. Its solution on partitioned domains requires specialized treatment due to its unique characteristics. Unlike the Poisson’s equation, the Helmholtz equation’s discretization typically produces a symmetric but non-positive definite matrix [14]. In the context of conventional numerical methods (e.g. finite difference method), separate domain decomposition methods have been proposed to tackle Poisson’s and Helmholtz equations (e.g. [29]). This is because transmission conditions that work well for the Poisson’s equation does not readily extend to handle the Helmholtz equation. To this end, solving the Helmholtz and Poisson’s equations with a unified domain decomposition framework presents substantial technical challenges.
Here, we adopt a complex-valued, 2D Helmholtz equation problem that has an exact solution. This problem was used by Babuška et al. [42] to validate a generalized finite element method for the Helmholtz equation to reduce the rising error with higher wave number. This example was also adopted in the work of Bao et al. [43] to validate a discrete singular convolution algorithm for solving the Helmholtz equation with high wavenumbers.
The exact solution to the problem is a planar wave of the form $u(x,y)=\exp{(j(k_{1}x+k_{2}y))}$, where $j$ is the imaginary unit, in a unit square $\Omega={(x,y)|0\leq x,y\leq 1}$ with a Sommerfeld-type radiation boundary condition as follows:
$-\Delta u-k^{2}u$ $=0,$ $\text{in}\quad\Omega,$ (21)
$jku+\frac{\partial u}{\partial n}$ $=g,$ $\text{on}\quad\partial\Omega,$
where $(k_{1},k_{2})=k(\cos\theta,\sin\theta)$ and the function $g$ depends on the wave direction $\theta$. Here, we set $k=\pi*2^{4}$ and $\theta=\dfrac{\pi}{3}$. We use a $4\times 4$ domain decomposition with a $128\times 128$ global uniform mesh. The neural network architecture for each subdomain is identical to that used in the previous Poisson’s equation cases, with the exception that it now produces two outputs, corresponding to the real and imaginary parts of the solution, respectively. In terms of the definition of boundary loss terms, the squared magnitude of a complex-valued residual is equivalent to the summation of the squared real part and imaginary parts. For instance, the complex value of the boundary residual operator $\mathcal{B}(\theta;\mathbf{x}_{i})$ on the $i$th boundary point is $\mathfrak{Re}[\mathcal{B}(\theta;\mathbf{x}_{i})]+j*\mathfrak{Im}[\mathcal{B}(\theta;\mathbf{x}_{i})]$, then the squared magnitude is
$||\mathcal{B}(\theta;\mathbf{x}_{i})||_{2}^{2}=\mathfrak{Re}^{2}[\mathcal{B}(\theta;\mathbf{x}_{i})]+\mathfrak{Im}^{2}[\mathcal{B}(\theta;\mathbf{x}_{i})].$ (22)
In the same way, we keep the same real-valued interface parameters $\mathbf{q}$ as in previous examples, and the interface loss term on the $i$th interface point can be written as
$\mathcal{J}(\theta;\mathbf{x}_{i})\approx\sum_{k=1}^{3}q_{k}^{2}||\mathcal{O}^{k}_{i}||_{2}^{2}=\sum_{k=1}^{3}q_{k}^{2}\mathfrak{Re}^{2}(\mathcal{O}^{k}_{i})+q_{k}^{2}\mathfrak{Im}^{2}(\mathcal{O}^{k}_{i}).$ (23)
Therefore, we separate the real and imaginary parts and avoid the occurrence of cross-product terms, as argued in Section 3.2, when implementing the complex-valued interface parameters. We note that our proposed method remains unchanged.
Table 2 summarizes the statistical performance of the proposed method over 5 independent trials, in terms of the relative $\mathit{l^{2}}$ and $\mathit{l^{\infty}}$ norms for real and imaginary parts of the solution. The mean and standard deviation of $\mathcal{E}_{r}$ with an order of magnitude of $10^{-4}$ suggest that an accurate solution can be predicted consistently with each trial. The levels of $\mathit{l^{\infty}}$ norm, which is a stricker norm than the $\mathit{l^{2}}$ as a metric of accuracy, further support this outcome.
Complex-valued prediction Relative $\mathit{l^{2}}$ norm $\mathit{l^{\infty}}$ norm
$\mathfrak{Re}$ part $43.76e-4\pm 4.87e-$ 4 $25.28e-2\pm 7.08e-$ 2
$\mathfrak{Im}$ part $41.46e-4\pm 1.28e-$ 4 $21.47e-2\pm 1.27e-$ 2
Table 2: The relative $\mathit{l^{2}}$ and $\mathit{l^{\infty}}$ norms for the real and imaginary parts of the predicted solution representing five independent trials.
Figure 7 shows the distribution of the predicted solution and point-wise error from the best trial. The wave propagation with a phase lag is visible on both real and imaginary parts of the solution in Fig. 7(a) and Fig. 7(c), respectively. The absence of wave reflection or distortion along the boundaries indicates that the Sommerfeld-type radiation boundary conditions are effectively handled by our PECANN method. From Fig. 7(b) and Fig. 7(d) we observe that the absolute error is evenly distributed across the domain, with no significant accumulation along the interfaces or boundaries.
Figure 7: 2D Helmholtz equation problem with a complex-valued solution: (a) real part of the predicted solution, (b) real part of the absolute point-wise error, (c) imaginary part of the predicted solution, (d) imaginary part of the absolute point-wise error from the best trial.
For ease of presentation, we select four subdomains out of the total 16. Figure 8 presents the evolution of the individual loss terms, relative $\mathit{l^{2}}$ norms and the two interface parameters chosen. From Fig. 8(a-d), we observe that both boundary and PDE constraints and the objective function as well as the relative error $\mathcal{E}_{r}$ rapidly decrease after around 1000 outer iterations. For outer iterations ranging from $1000$ to $1600$, interface parameters $\alpha$ and $\beta$ shown in Fig. 8(e) and Fig. 8(f), respectively, plateau at a value around 50. After $1600$ outer iterations, the interface parameters continue their upward trend; however, the objective functions converge to low values, indicating that the interface operators are being satisfactorily met.
Figure 8: Helmholtz equation with a complex-valued solution: Evolution of (a) PDE constraints, (b) boundary condition constraints, (c) objective functions, (d) relative $\mathit{l^{2}}$ norms, (e) interface parameter $\alpha$, (f) interface parameter $\beta$ for four randomly-selected subdomains.
4.1.4 Helmholtz equation with a real-valued, high-wavenumber solution
@@
We use the same uniform global mesh of $160\times 160$ as adopted in [25]. We partitioned our domain with a $5\times 5$ decomposition. In our Algorithm 1, the penalty scaling factor for boundary condition constraints, $\eta_{B}$, is set to 10, and the smoothing constant $\zeta$ is set to 0.999. Our neural network and optimizer configurations remain identical to those we have used in the Poisson’s equation investigations. The exact solution to this problem is not known. Therefore, to evaluate the quality of the predictions, the same finite difference method, as adopted in [25], on a mesh of $360\times 360$ is used to obtain the reference solution.
Figure 9: Helmholtz equation with a real-valued, high-wavenumber solution: (a) finite difference solution on a $5\times 5$ partitioned domain, (b) predicted solution, (c) absolute point-wise error.
@@
Figure 10: Helmholtz equation with a real-valued, high-wavenumber solution: Evolution of (a) PDE constraints, (b) boundary condition constraints, (c) objective functions, (d) relative $\mathit{l^{2}}$ norms, (e) the ratio $\beta/\alpha$, (e) the ratio $\gamma/\alpha$ for six specific subdomains.
To illustrate the performance of the method, we select six specific subdomains based on the symmetry of the solution, and track the evolution of the loss terms, the relative $\mathit{l^{2}}$ norms, and the interface parameter ratios, as shown in Figure 10. At the early stages of training, most subdomains, except the central one (2,2), fall into local minima with low values for PDE, boundary constraints and objective functions, as shown in Figures 10(a-c). This behavior arises from the challenging problem setup: the source term, with its narrow width and high magnitude, is concentrated in the central subdomain, while the surrounding subdomains have near-zero source terms. The high wave number further complicates the reduction of the PDE constraint in the central subdomain. As the PDE constraint in this region decreases to around $10^{-3}$, its solution improves, gradually propagating information to neighboring subdomains. This leads to an increase in the objective functions (interface losses) in adjacent subdomains, pulling them out of their local minima, as reflected in the noticeable rise in constraint values. The interplay between satisfying the boundary conditions and improving the physics constraints drives the reduction of the relative $\mathit{l^{2}}$ norms, as seen in Figure 10(d). This process continues over 30,000 iterations towards convergence of $\mathcal{E}_{r}$.
@@ 4.2.1 Inferring thermal conductivity in multilayered materials
The technical challenge lies in inferring two independent $\kappa$ values from sparse data. We generate synthetic measurement data at ten locations using pseudo sensors that are equidistantly positioned. These sensors are placed between $y=0.2$ and $y=0.8$, situated 0.05 units away from the interface. Using a $2\times 1$ domain decomposition, each subdomain is assigned to a distinct layer. The global domain includes $64^{2}$ randomly distributed collocation points, comprising residual, boundary, and interface points, as shown in Figure 11. For evaluation, we employ a uniform $128\times 128$ mesh. The same neural network architecture, hyperparameters, and optimizer settings from the forward Poisson’s problem, as shown in Figure 2, are used here. It is important to note that the additional data constraint shares the same default penalty scaling factor, $\eta_{B}=1$, as the boundary condition constraint. The unknown conductivity of each layer is treated as a model parameter to be learned, i.e., $\kappa_{k}\in\theta_{k}$ for $k\in\{1,2\}$, with initial values randomly assigned between 0 and 1.
@@
Figure 13 depicts the evolution of two accuracy metrics, constraints, objective functions, and interface parameters, as well as their ratios for both layers. The relative $\mathit{l^{2}}$ norms of the temperature and the absolute errors of two predicted conductivities, $\mathit{l^{\infty}}$, are shown in Figs. 13(a-b), respectively. In the beginning, both metrics exhibit rapid convergence, with the errors in $\mathcal{E}_{r}^{u}$ and $\mathcal{E}_{\infty}^{\kappa}$, separately decreasing by approximately five and three orders of magnitude. After a brief stabilization phase, further reductions occur around 700 outer iterations, coinciding with a slight decrease in the boundary condition and data constraints, in Fig. 13(c), and a notable rise in the objective functions, in Fig. 13(d). The increase in objective function values is attributed to the growth of interface parameters, seen in Fig. 13(e). As discussed in Section 3.2, increasing $q^{j}$ helps maintains a strong gradient when the corresponding operator $\mathcal{O}_{i}^{j}$ converges to small values, thereby enhancing convergence rate. Eventually, the objective functions stabilize, and both accuracy metrics reach convergence after approximately 1200 outer iterations. Additionally, Fig. 13(f) shows that the interface ratios also stabilize between 1 and 2.
@@ 4.2.2 Inferring thermal conductivity in functionally graded materials
Here, we adopt a similar problem used by Shukla et al. [23]. Instead of the map of the United States of America, we consider a butterfly-shaped domain $\Omega$. The geometry of the domain is defined as:
@@
Figure 15 displays the predicted distributions of temperature and thermal conductivity, along with their absolute errors. In Fig. 15(b), the temperature prediction errors are relatively uniform across the four subdomains, generally within the order of $10^{-3}$ and are primarily localized near the boundaries. In contrast, Fig. 15(d) reveals that the errors in conductivity predictions are mostly concentrated in the top subdomains, particularly in $(0,0)$, where the maximum error is $1.033\times 10^{-1}$. Comparing these errors with the true ranges shown in Figs. 15(a) and 15(c), both predictions demonstrate high accuracy, with temperature predictions being more precise than those for conductivity.
Figure 16: Inverse identification of spatially varying conductivity in functionally graded materials: Evolution of (a) relative $\mathit{l^{2}}$ norms for temperature, (b) relative $\mathit{l^{2}}$ norms for conductivity, (c) objective functions, (d) PDE constraints, (e) boundary condition constraints, (f) data constraints, (g) interface parameter ratio $\alpha^{(k)}/\alpha^{(u)}$, (h) $\beta/\alpha^{(u)}$, (i) $\gamma/\alpha^{(u)}$ for all the four subdomains.
Figure 16 shows the evolution of the relative $\mathit{l^{2}}$ norms for temperature and conductivity, the objective functions, three constraints, and the interface parameter ratios across all subdomains. As seen in Fig. 16(a), $\mathcal{E}_{r}^{u}$ converges first, oscillating around $10^{-4}$ after approximately 2500 outer iterations, while $\mathcal{E}_{r}^{\kappa}$, in Fig. 16(b), continues to decrease gradually. Around the same outer iterations, the objective functions, which initially rise, begin to decrease rapidly in Fig. 16(c). The PDE, BC and data constraints, in Figs. 16(d-f), reduce to approximately $10^{-6}$, $10^{-6}$, and $10^{-5}$, respectively. Among these, subdomain (1,0) exhibits a slower decay in $\mathcal{C}_{B}$, but ultimately achieves the lowest $\mathcal{E}_{r}^{\kappa}$ over the iterations. Due to the inclusion of the Dirichlet interface operator for $\kappa$, Fig. 16(g) shows the ratios of the interface parameters $\alpha^{(\kappa)}$ to $\alpha^{(u)}$. $\dfrac{\alpha^{(\kappa)}}{\alpha^{(u)}}$, as well as the other interface parameter ratios, $\dfrac{\beta}{\alpha^{(u)}}$ and $\dfrac{\gamma}{\alpha^{(u)}}$ in Figs. 16(h-i), fluctuate between 1.2 and 1.5 after an initial spike. As noted earlier, the convergence of the relative $\mathit{l^{2}}$ norms does not necessarily require the convergence of the interface parameter ratios.
@@ 4.3 Parallel Performance
We assess the parallel efficiency of our domain decomposition method, using the same Poisson’s equation problem as outlined in Section 4.1.1. Using the Adam optimizer, we evaluate both weak and strong scaling performances, varying the number of subdomains from 2 to 32. The performance numbers are derived from the average time costs recorded over 5 independent trials for each configuration. The numerical experiments are conducted on an Intel Xeon Platinum 8462Y+ architecture. Each subdomain is managed by an MPI process (mpi4py 3.1.5 package) bound to a single CPU core. A Cartesian communicator is created via Create_cart function and the interface information is sent and received via Isend and Irecv functions, respectively.
Our parallel implementation is straightforward, and we have not explored more advanced techniques, such as overlapping communication with computation, to improve scalability. Additionally, our performance analysis is limited to a maximum of 32 processes. It is well known that DDM typically do not scale efficiently to a large number of subdomains without a coarse space correction [15], which has not been pursued in this study. These limitations should be taken into account when judging the scalability of our approach based on the following results.
@@ 4.3.1 Weak Scaling Analysis
Figure 17(a) captures the runtime for five outer iterations, segmented into computation and communication times. The total runtime slightly increase, approximately 60 seconds, across various domain decompositions, with computation time significantly exceeding communication time. This predominance is attributed to the design of our domain decomposition training procedure, in which communications between neighboring subdomains are scheduled for after at least 50 inner epochs. This is possible because our PECANN framework aims to infer the parameters of a generalized interface condition constrained by the PDE and its boundary conditions, as opposed to learning the parameters of a neural network constrained by an uncertain interface transmission condition.
@@ 5 Conclusion
Domain decomposition methods (DDMs) are crucial for advancing physics-informed or constrained machine learning techniques to solve large-scale partial differential equations (PDEs) in a distributed setting. We introduce a novel Schwarz-type, non-overlapping DDM to learn the solution of forward and inverse PDE problems using physics-informed neural networks. Specifically, we leverage the physics and equality-constrained artificial neural networks (PECANN) framework [26] in a novel way, formulating a constrained optimization problem for each subdomain. An augmented Lagrangian method with a conditionally adaptive strategy, unique to this study, is employed to transform each constrained optimization problem into a dual unconstrained problem.
Unlike the original PECANN method, which uses boundary conditions exclusively to constrain the PDE loss term, our DDM allows both the boundary conditions and governing PDE to jointly constrain a generalized interface loss term within the PECANN framework. This formulation also differs from other physics-informed neural network approaches with DDM, where subdomain interface conditions are used alongside boundary conditions to softly constrain the PDE solution.
Our proposed interface transmission condition is formulated as a linear combination of Dirichlet, Neumann, and tangential derivative continuity operators, each with subdomain-specific interface parameters to be optimized. Crucially, we derive an approximate interface loss term based on the mean squared error estimator in which the cross product terms of the interface operators are neglected, resulting in a formulation (i.e. Eq. 13) more suitable for optimization. The PDE solution is then learned via a feed-forward neural network while minimizing this approximate interface loss term simultaneously. Subdomains are independently trained using separate neural networks with the same architecture for a fixed number of epochs before synchronizing with neighboring subdomains, thus minimizing communication overhead while promoting the independent learning of interface parameters. We find that the augmented Lagrangian method with the proposed conditionally adaptive strategy plays a crucial role in the optimization process. It is also worth noting that our proposed interface transmission condition may not be directly applicable as a DDM for classical numerical methods, such as finite differences, due to the fundamentally different approach to predicting the solution. In our approach, the solution is learned by minimizing the approximate interface loss term, rather than by directly enforcing the interface transmission conditions.
We have applied our method to learn the solution of a variety of forward and inverse PDE problems in parallel with up to 64 subdomains using a Message Passing Interface model. A promising aspect of our method, facilitated by the proposed approximate interface loss term, is its capability to learn solutions to both Poisson’s and Helmholtz equations, including those with high wavenumber as well as complex-valued solutions. This unified framework for tackling PDEs with fundamentally different characteristics stands in sharp contrast to the state-of-the-art in conventional DDMs, which usually require distinct interface transmission conditions for equations like Poisson’s and Helmholtz. It is also important to note that, despite the common underlying framework, the objective functions remain specific to each PDE due to the PDE-constrained nature of minimizing a subdomain-specific interface loss term. Additional numerical experiments would be valuable for thoroughly evaluating the potential of our approach in solving PDEs with varied characteristics and increasing complexity.
Our method exhibits consistent generalization capabilities as the number of subdomains increases; however, this trend is inherently limited and cannot continue indefinitely. Our parallel performance analysis has so far been limited to 32 processes, and further testing is needed to assess parallel scalability beyond this number. It is well-documented that Schwarz-type domain decomposition methods often experience scalability challenges when applied across a large number of subdomains [15], which we expect to be the case for our method as well. In such cases, introducing a coarse space or level correction is recommended to improve scalability of parallel solutions with numerous subdomains. A promising strategy to address this challenge could be the approach proposed by Mercier et al. [57] in the context of deep DDM [58]. In our approach, a coarse level can be introduced by representing the solution with a neural network that spans the entire domain. Each subdomain can communicate information sampled from its interior points to the coarse-space neural network in a data-driven manner, which helps to constrain the solution at the coarse level. The augmented Lagrangian method within our PECANN framework is particularly well-suited for assimilating such data into the learning process. When solving within each subdomain, the generalized interface conditions can be enriched by incorporating information from the global coarse-space solution. This information can be weighted appropriately, with the weight being learned alongside other subdomain-specific interface parameters.
It is essential to highlight the limitations of the proposed method. Solving PDEs using neural networks is significantly more time-consuming than conventional numerical methods, particularly for forward problems. This time burden becomes even greater for complex PDEs with multi-scale or high-wavenumber features in their solutions, where prolonged training times are necessary to achieve accurate predictions. Furthermore, neural network-based methods often require multiple runs to ensure solution accuracy, further increasing computational demands. Although we observe that increasing the number of subdomains enhances the accuracy of learned solutions due to our method’s generalization capabilities, this improvement has inherent limits. A coarse-space correction could potentially address this limitation, which we plan to explore in future work, along with an in-depth scalability assessment involving a larger number of subdomains and applications to complex three-dimensional problems.